# UVLM: Benchmarking Video Language Model for Underwater World Understanding

### v1

@@ 3.4 Evaluation Metric
Inspired by LLM judge [55] and MMDU [56], we build the evaluation pipeline by using GPT-4o. Specifically, based on GPT-4o, we evaluate the predictions from six aspects: 1) Semantic accuracy, assesses how closely the description matches the reference answer; 2) Visual perception accuracy, examines the correctness of the model’s interpretation of image content; 3) Detail completeness, measures the thoroughness of the provided information; 4) Environmental description accuracy, focuses on the correctness of underwater environmental descriptions; 5) Species behavior matching, evaluates the accuracy of prediction; 6) Choice accuracy, assesses the correctness of multiple-choice question. Please refer to the Appendix B for our evaluation prompts.
4 Experiments
Table 3: Performance comparison of different methods.
Method
Multiple
Choice
@@ 4 Experiments
Closed-source VidLMs and other Open Source Large VidLMs
GPT4o 77.72 77.67 73.40 78.23 80.07 79.73 77.95
Claude3.7-Sonnet 76.61 73.35 73.58 74.10 79.71 76.35 76.09
Gemini2.5-Flash 78.22 72.43 73.34 70.53 78.32 74.92 75.00
Qwen2.5VL-72B 75.97 74.22 71.94 74.85 78.45 77.40 75.49
Base VidLMs
InternVL2.5-1B 48.54 44.35 47.30 45.30 50.15 44.85 46.73
VideoLLaMA3-2B 57.17 55.89 58.48 58.53 63.15 55.74 58.39
InternVL2.5-8B 57.64 57.70 59.75 59.45 63.48 61.05 60.15
VideoLLaMA3-7B 63.83 60.43 62.07 62.05 66.97 61.41 62.70
InternVL2.5-1B + UVLM 64.06 54.63 57.14 55.52 64.06 59.14 59.04
+15.52 +10.28 +9.76 +10.22 +13.91 +14.29 +12.31
VideoLLaMA3-2B + UVLM 71.96 63.66 65.35 65.74 72.48 67.70 67.61
+14.79 +7.77 +6.87 +7.21 +9.33 +11.96 +9.22
InternVL2.5-8B + UVLM 70.92 65.55 66.10 66.55 70.45 70.17 68.25
+13.28 +7.85 +6.35 +7.10 +6.97 +9.12 +8.10
VideoLLaMA3-7B + UVLM 76.92 70.52 70.17 71.39 76.41 73.39 73.03
+13.09 +10.09 +8.10 +9.34 +9.44 +11.98 +10.33
4.1 Dataset Partition
The proposed UVLM contains 2109 videos, covering 419 categories. As shown in Figure 3 (b), the overall category distribution follows a long-tail distribution and videos of many specific categories have relatively few observed targets. To ensure the training and test sets have consistent category distributions as much as possible, we sampled the test set proportionally. For categories with more than 5 videos, we randomly selected videos for the test set according to a predefined train-test split ratio. Then, we further randomly sampled videos from categories with fewer than 5 videos as test samples. Finally, 208 videos and their corresponding video-text pairs were selected as test set.
4.2 Results on UVLM
Table 3 and Figure 4 present the results on UVLM. We selected two recently released VidLMs, InternVL and Video-LLaMA, as baselines. For InternVL, we evaluated the 1B and 8B variants, while for Video-LLaMA, we tested both the 2B and 7B versions. According to the experimental results, we can draw several observations. First, underwater observation is indeed different from in-air scenarios. Even for current SOTA closed-source models like GPT-4o and Gemini, the scores only reach 77.95 and 75.00, respectively. Among open-source models, Qwen2.5VL-72B achieves a score of just 75.49. This indicates that there remains significant room for further improvement in underwater world understanding capabilities. Second, fine-tuning with UVLM significantly enhances VidLMs’ underwater observation capabilities. For instance, VideoLLaMA3-7B achieves a more than 10 points improvement in accuracy after fine-tuning, reaching an overall score of 73.03. At just 7B parameters, it trails Qwen2.5VL-72B (75.49) by only 2.46 percentage points, which demonstrates remarkable efficiency. Moreover, its performance becomes comparable to closed-source models like Gemini (75.00), despite the latter’s likely larger scale and proprietary training data. In fact, the benefits of UVLM fine-tuning extend beyond underwater observation, it also enhances performance on traditional benchmarks such as VideoMME and Perception Test. For detailed experimental results, please refer to the Appendix C.
5 Conclusion
In this paper, we introduced the first multimodal underwater observation benchmark specifically designed for enhancing video language model applications within underwater environments. Our benchmark integrates diverse and rich datasets, combining human expertise and advanced AI annotation methods to create an extensive resource suitable for multimodal instruction tuning. This dataset uniquely includes videos with varied resolutions, frame rates, and extensive marine biodiversity, covering 419 marine animal classes along with diverse static plants and terrains. Additionally, we meticulously addressed inherent underwater challenges, such as varying illumination, turbidity, and diverse viewing angles, providing realistic conditions for robust model training and evaluation. The benchmark incorporates comprehensive task diversity across 20 distinct observation scenarios, ranging from static content and dynamic behaviors to static and dynamic scene analyses. By proposing carefully designed and challenging evaluation metrics, our benchmark enables meaningful quantitative comparisons and facilitates the development of high-accuracy, reliable underwater observation systems. We anticipate this dataset will serve as a foundational resource for future research, encouraging innovation and promoting environmentally sustainable and economically viable solutions for underwater monitoring and exploration.
Figure 4: Visualization examples of fine-tuning with UVLM. Error descriptions are marked in red, and the accurate descriptions are marked in green.
6 Limitation & Broader Impacts
UVLM is the first large-scale underwater video–language benchmark, spanning 419 marine species, 20 task types. This rich coverage provides a solid foundation for robust, fair comparison of multimodal models in challenging aquatic conditions. While a small fraction of annotations may contain noise or class imbalance and extreme scenarios warrant further expansion, the present scale and quality already enable reliable training and analysis. Expert cross-checks, detailed documentation, and clear usage terms mitigate potential bias and misuse, while inviting the community to broaden geographic and modality coverage for improved generalization. UVLM is poised to accelerate marine-ecosystem monitoring, sustainable fisheries management, and public education, fostering cross-disciplinary collaboration and helping establish a secure, transparent, and responsible AI ecosystem for underwater research.

### v2

@@ Introduction
UVLM’s comprehensive evaluation employs two Objective Metrics and five LLM-based Judgement Metrics (details in the Appendix.C). Two objective metrics are: 1) Multiple Choice Accuracy, evaluate the model’s performance on various common multiple-choice questions; 2) Fine-grained taxonomic classification, evaluate the model’s performance on multiple-choice questions related to the professional biological taxonomy (Whittaker 1969). The LLM-based Judgement Metrics are inspired by LLM judge (Zheng et al. 2023) and MMDU (Liu et al. 2024c), leveraging GPT-4o as the evaluation backbone. Specifically, they assess model performance from five aspects; 1) Semantic accuracy, assesses how closely the description matches the reference answer; 2) Visual perception accuracy, examines the correctness of the model’s interpretation of image content; 3) Detail completeness, measures the thoroughness of the provided information; 4) Environmental description accuracy, focuses on the correctness of underwater environmental descriptions; 5) Species behavior matching, evaluates the accuracy of prediction. Evaluation prompts are available in Appendix.C.
Experiments
Dataset Partition
UVLM contains 2109 videos, covering 419 categories. As shown in Figure 3 (b), the overall category distribution follows a long-tail distribution and videos of many specific categories have relatively few observed targets. To ensure the training and test sets have consistent category distributions as much as possible, we sampled the test set proportionally. For categories with more than 5 videos, we randomly selected videos for the test set according to a predefined train-test split ratio. Then, we further randomly sampled videos from categories with fewer than 5 videos as test samples. Finally, 208 videos-text pairs were selected as test set.
Experimental Results on UVLM
Table 2 and Figure 4 present the results on UVLM. We selected three recently released VidLMs (InternVL (Chen et al. 2024), QwenVL (Bai et al. 2025), VideoLLaMA (Zhang et al. 2025a)) and a representative VLM (Qvis (Lu et al. 2024)) as baselines. For InternVL, we evaluated the 1B and 8B variants, while for VideoLLaMA, we tested both the 2B and 7B versions. According to the experimental results, we can draw several observations:
  • Underwater observation poses distinct challenges compared to in-air scenarios. Even state-of-the-art closed-source models, GPT-4o (77.95) and Gemini (75.00), or open-source models like Qwen2.5VL-72B (75.49), achieve relatively limited performance. The gap highlights unique complexities of underwater scenarios, suggesting significant potential for further exploration
  • Fine-tuning with UVLM significantly enhances VidLMs’ underwater observation capabilities. For example, VideoLLaMA3-7B achieves an accuracy gain exceeding 10 points, reaching 73.04, just 2.45 points behind the much larger Qwen2.5VL-72B (75.49). Remarkably, this compact model also closely matches closed-source models like Gemini (75.00), despite their larger scales and proprietary datasets. These results indicate that the proposed human-AI collaborative annotation pipeline effectively distills knowledge from large models like GPT-4o and injects it into smaller models through fine-tuning. While maintaining acceptable performance trade-offs,it significantly reduces model overhead and hardware requirements and expand the model’s applicability scope.
  • In highly specialized fields, simply increasing the amount of training data is insufficient to bridge the gap between small models and large models. Differing from other six metrics, performance on the fine-grained taxonomic classification task requires complex biological domain knowledge. For this particular task, we observe a persistent performance gap between small and large models even after extensive fine-tuning. We attribute this phenomenon to the fundamental capacity differences between model scales. While simple QA tasks can be effectively addressed through data fine-tuning alone as they impose relatively low demands on model capacity, tasks requiring sophisticated domain expertise present a greater challenge. Small models, constrained by their limited capacity while simultaneously maintaining performance across multiple tasks, demonstrate particular difficulty in closing this performance gap through fine-tuning alone. This finding presents a new challenge for the field of underwater world understanding.
In addition, the benefits of UVLM fine-tuning extend beyond underwater observation, improving performance on general benchmarks such as VideoMME and Perception Test. Additional results are provided in the Appendix.E.
Conclusion
This paper introduces the first multimodal benchmark specifically designed to enhance video-language model applications in underwater environments. Combining human expertise with advanced AI annotation methods, UVLM encompasses extensive marine biodiversity (419 animal classes), diverse video resolutions, and realistic underwater challenges, including varying illumination and water turbidity, thus providing authentic conditions for model training and evaluation. Featuring 20 comprehensive subtasks across distinct scenarios with carefully designed evaluation metrics, UVLM enables meaningful quantitative comparisons, fostering the development of accurate and reliable underwater observation systems. We anticipate this resource will promote innovation and contribute to sustainable underwater monitoring and exploration.
Appendix A Appendix Contents & Benchmark Link
  • A. Sample Generate Prompt Details & Annotation Error Statistics
  • B. Task Structure & Example Visualization of UVLM
  • C. Objective Metrics & LLM-based Evaluation Prompt Details
  • D. Finetuning Setting Details
  • E. Evaluation on Generalization Effect of UVLM Fine-tuning
  • F. Evaluation with Different Judgment Models
  • G. Evaluating How Underwater Image Quality Influences Model Accuracy
Figure 1: Dialogue generation prompt
Appendix B A. Sample Generate Prompt Details
In this section, we provide the prompts we used to generate the samples, as shown in Figure 1.
First, our prompt specifies the domain to which the video belongs and the coordinate location of the target object, guiding GPT-4o to pose insightful and relevant questions aligned with the visual theme. We leverage GPT-4o to generate a large number of questions centered on the underwater world, ensuring that these questions comprehensively cover the 20 thematic domains we have proposed. Secondly, to ensure fairness in data generation, we do not provide any textual cues when generating answers. Instead, GPT-4o is required to understand and respond to multiple questions based solely on the video content.
In the initial part of the prompt, as shown in Figure 1, we assign GPT-4o the role of a marine video analysis expert, explicitly framing the task as designing and answering diverse multi-turn questions for a video featuring marine life. This role assignment encourages the model to draw from disciplinary frameworks such as marine biology, ecology, and behavioral science, ensuring that the output remains coherent and scientifically grounded. Immediately following this, the prompt provides an overview of the reference materials, including annotated video frames and supplemental cues. These materials serve as core visual and contextual anchors, implicitly guiding the reasoning process behind the generated questions and answers.
The main body of the prompt outlines the key requirements that the model must follow when generating questions and answers, ensuring comprehensive coverage of marine organism research. These requirements are structured into several subfields:
  • Biological Characteristics: The model is expected to observe and infer the organism’s body size, coloration, habitat, dietary habits, and social behaviors, as well as possible predation or foraging strategies.
  • Observed Target Characteristics: This focuses on the organism’s specific behaviors, swimming posture, action sequences, and any signs of health or emotional state.
  • Illustrated Environmental Features: The model should describe the environmental conditions depicted in the video, such as substrate composition, water clarity, and lighting.
  • Physicochemical Water Properties: This includes attention to water chemistry and physical parameters (e.g., salinity, temperature, current strength), allowing for deeper inference into species adaptability and ecological preferences.
These structured guidelines require the model to integrate information across biological, ecological, and environmental domains, resulting in questions and answers with greater scientific depth and contextual richness.
Figure 2: Task structure and the examples of UVLM. Left side: Task structure; Right side: several cases
Method Acc_obj (%) Acc_subj (%)
GPT-4o 45 47
+PK 91 95
+PK+MC 100 100
Table 1: Annotation accuracy on objective and subjective labels under different settings. PK: prior knowledge; MC: manual correction; Acc_obj/Acc_subj (%): annotation accuracy on objective/subjective labels.
Annotation statistical reliability: As summarized in Table 1, incorporating prior knowledge (+PK) boosts GPT-4o from 45%/47% to 91%/95% accuracy on objective/subjective labels, indicating that most errors stem from missing domain priors rather than random noise. Adding manual correction (+PK+MC) further removes all residual errors on the curated set, yielding 100% accuracy for both label types. During correction, 9.5% of videos were discarded due to irreconcilable ambiguity or low visual quality; thus, the true accuracy on the original pool is conservatively bounded between 90.5% and 100%. Inter-expert agreement on the remaining items ranges from 91.67% to 100% for two-expert groups and from 78% to 100% for three-expert groups, demonstrating high annotation consistency.
Appendix C B. Task Structure & Examples of UVLM
The task structure and several cases of UVLM are shown in the left side and right side of Figure 2 respectively. In UVLM case 1, a fish in the upper-left corner of the frame becomes the focal point of attention. By observing its steady swimming path and the surrounding environment, it can be inferred that the fish is slowly moving deeper into the kelp forest area (Option B), demonstrating a high degree of environmental adaptation and a clear exploratory intent. Throughout its journey, the fish remains within the cover of the kelp, avoiding open waters and showing no significant changes in depth. This cautious movement strongly suggests that its behavioral pattern aligns with ”navigating cautiously through dense kelp” (Option A).
Visually, the fish features a silvery body tone with a slightly shadowed dorsal area. This coloration provides natural camouflage in the dim kelp forest, helping reduce predator detection and enhancing foraging efficiency. The surrounding environment appears stable and characteristic of a typical kelp forest habitat, eco-complex and suitable for the survival of various small organisms. Based on this, it is inferred that the fish is omnivorous, primarily feeding on small invertebrates and algae, which is a common dietary structure within kelp forest ecosystems. Taken together, the four interrelated questions not only reveal the fish’s movement direction and behavioral patterns but also connect its physiological traits with its ecological adaptation strategies. This makes the sequence a representative example of fish behavioral ecology within a kelp forest environment.
Figure 3: Judgment prompt.
Figure 4: Several evaluative demos of UVLM.
Table 2: Representative video language benchmarks
Name Venue Themes & Scenarios
MSVD-QA (Xu et al. 2017), Next-QA (Xiao et al. 2021) $M\prime$17, CVPR $\prime$21 Daily life
[…]
EgoSchema (Mangalam, Akshulakov, and Malik 2023) NeurIPS $\prime$23 Life and human action, egocentric perspective
Perception Test (Patraucean et al. 2023) NeurIPS $\prime$23 Indoor scenes (living room/kitchen), with a small number in bathroom or outdoor settings
Video-Bench (Ning et al. 2023) arXiv $\prime$23 Television, music videos, NBA highlights
InternVid (Wang et al. 2023) ICLR $\prime$23 People and blogs, education, news and politics, how-to and style, entertainment, gaming, science, sports
MovieChat-1K (Song et al. 2024) CVPR $\prime$24 Movie clips
MVBench (Li et al. 2024) CVPR $\prime$24 Life, human action, movies, people, pets, science, instruction, finance, news, sports, humor, etc.
LONGVIDEOBENCH (Wu et al. 2024) NeurIPS $\prime$24 Life, movie, knowledge and news
E.T. Bench (Liu et al. 2024a) NeurIPS $\prime$24 Indoor activities, tabletop, sports, egocentric views, cooking, news and vlogs, how-to, open-ended
LVBench (Liu et al. 2024b) arXiv $\prime$24 Sports documentary, event record, lifestyle, TV show, cartoon
EgoThink (Cheng et al. 2024) CVPR $\prime$24 Daily life understanding from first-person perspective
[…]
Appendix D C. Objective Metrics & LLM-based Evaluation Prompt Details
In Figure 3, we present the evaluation prompt used to direct GPT-4o in rigorously scoring model’s responses against ground-truth answers. The assessment framework evaluates performance across seven key dimensions, including two objective metrics (multiple-choice accuracy and fine-grained classification) along with five LLM-based judgment metrics (semantic accuracy, detail completeness, visual perception accuracy, environmental description accuracy, and species behavior matching). Multiple-choice accuracy is determined by converting performance on four single-answer questions into a percentage-based score. The fine-grained classification component specifically assesses knowledge of aquatic organism phylogenetics through multiple-choice questions covering four taxonomic levels (kingdom, phylum, class, and order), with strict scoring rules: each omitted correct option deducts 40 points (one omission) or 80 points (two omissions), while any incorrect selection results in an automatic zero for that question. For these two metrics, the LLM serves solely as a statistical tool to measure the discrepancy between model responses and ground truth, ensuring the evaluation is objectively calculated. For the five LLM-based metric, each metric is scored on a 10 to 100 scale using detailed evaluation criteria. For the five LLM-based metrics, performance is measured across five progressive scoring tiers (10-20, 20-40, 40-60, 60-80, and 80-100) with increasingly stringent requirements. This comprehensive scoring system ensures rigorous evaluation across all assessment dimensions. The judge must justify every sub-score with concise evidence-based reasoning, then derive an Overall Score that synthesises the dimension-level results.
To clearly illustrate the testing and evaluation of UVLM, we present several Q&A pairs from UVLM in Figure 4 Due to space limitations, we are unable to show complete conversations; therefore, each case includes only two representative Q&A pairs: one multiple-choice question and one open-ended question. Each case is accompanied by the corresponding video and the ground truth answers. In addition, we list the results of both VideoLLaMA3-7B and VideoLLaMA3-7B+UVLM. We also provide evaluation scores generated by GPT-4o, including the reasoning behind the scores and the specific ratings.
Appendix E D. Finetuning Setting Details
In the experimental section, we finetuned the InternVL2.5 and VideoLLaMA3 models using our self-constructed UVLM dataset. The detailed evaluation metrics are presented in Table 3. It summarizes the key hyperparameter configurations used during the training of the two multimodal models: VideoLLaMA3 and InternVL2.5. The two models share most of the settings, including a batch size of 128, cosine learning rate scheduling, a warm-up ratio of 0.03, the AdamW optimizer, numerical precision of bfloat16, and the distributed training framework DeepSpeed ZeRO-3. However, InternVL2.5 differs from VideoLLaMA3 in two notable aspects: it adopts a smaller learning rate and applies a non-zero weight decay.
Table 3: Training hyper-parameters
Hyper-parameter VideoLLaMA3 InternVL2.5
batch size 128 128
learning rate 1e-5 2e-6
learning rate schedule cosine cosine
warm-up ratio 0.03 0.03
weight decay 0.0 0.01
epoch 1 1
optimizer AdamW AdamW
float precision bfloat16 bfloat16
deepspeed configuration zero3 zero3
Appendix F E. Evaluation on Generalization Effect of UVLM Fine-tuning
Motivation
Underwater data exhibit a pronounced domain gap relative to conventional terrestrial scenario benchmarks, and training on them directly may introduce capability bias. Fine-tuning on UVLM lets us ask a simple but powerful question: Does exposure to a visually and physically distinct domain (underwater) teach a general VLM something that transfers back to everyday terrestrial videos?
VidLMs are commonly trained on terrestrial footage that implicitly encodes strong priors: gravity points downward, illumination is quasi-static, and hues are only mildly distorted. Underwater scenes violate each of these assumptions. Severe wavelength-dependent absorption, dynamic caustics, six-DoF camera drift, buoyancy-driven motion, and turbidity collectively create an extreme domain shift. We hypothesize that forcing a VidLM to master these conditions regularizes its spatio-temporal representations, yielding better out-of-domain performance. To test this, we fine-tune a general VLM on our underwater dataset, UVLM, and evaluate on three terrestrial benchmarks.
Experimental Terrestrial Benchmarks
Table 2 provides a concise summary of representative benchmarks, highlighting their respective content coverage. Although these efforts have significantly advanced video-language research in terrestrial scenario contexts, the underwater environment remains largely uncharted, motivating our investigation of underwater video-language benchmark. In the following, we provide brief introductions to several well-established video-language benchmarks.
MVBench is a comprehensive benchmark designed to evaluate the performance of large multimodal language models on complex video understanding tasks. It consists of 20 diverse tasks that cover various aspects of video comprehension, including action recognition, object interaction detection, scene change detection, and fine-grained action classification. The benchmark challenges models to process both spatial and temporal information, testing their ability to reason about dynamic content in videos.
Figure 5: Several demos of comparison on related benchmarks.
The MVBench dataset includes 500 videos from various domains, each accompanied by over 10,000 multiple-choice questions and answers. These questions are designed to assess models’ understanding of key elements in video content, such as object presence, action sequences, and contextual reasoning over time. The dataset also includes questions that require models to perform more complex reasoning, such as predicting future actions or understanding interactions between objects within a scene.
Video-MME is a comprehensive multimodal benchmark dataset designed to evaluate model performance on complex video understanding tasks. It has over 900 video clips and 2,700 question-answer pairs. Each video is accompanied by audio and textual annotations, covering a wide range of real-world scenarios. The dataset spans six diverse visual domains: knowledge, film and television, sports competitions, artistic performances, daily life recordings, and multilingual content. These videos also vary significantly in duration, posing challenges for models to process and integrate visual, auditory, and textual information. The benchmark is intended to assess models’ ability to recognize fine-grained details, actions, and to understand contextual nuances across modalities.
Perception Test is a multimodal video benchmark designed to evaluate the perceptual and reasoning capabilities of pretrained multimodal models. It focuses on a range of perception skills, including memory, abstraction, physical understanding, and semantics, across video, audio, and text modalities. In addition, it evaluates various types of reasoning, such as descriptive, explanatory, predictive, and counterfactual reasoning. The dataset comprises 11,600 real-world videos and includes six types of annotations, such as multiple-choice questions, video-based Q&A, object and point tracking, temporal actions, and audio clips.
Results and Discussion
Table 4: Result on other benchmark
Method MvBench Perception Test VideoMME
InternVL2.5-1B 63.2 59.8 47.7
InternVL2.5-1B + UVLM 63.3 59.6 47.9
+0.1 -0.2 +0.2
VideoLLaMA3-2B 65.7 68.0 59.4
VideoLLaMA3-2B + UVLM 65.9 68.2 59.2
+0.2 +0.2 -0.2
InternVL2.5-8B 72.1 68.4 59.7
InternVL2.5-8B + UVLM 72.3 68.3 60.3
+0.2 -0.1 +0.5
VideoLLaMA3-7B 70.0 73.0 65.6
VideoLLaMA3-7B + UVLM 70.1 73.1 65.8
+0.1 -0.1 +0.2
Table 5: Evaluate the results of the four models using different judgment models.
Judgement
Model
Model
Semantic
@@ Appendix A Appendix Contents & Benchmark Link
GPT4o Claude-3.7-Sonnet 73.35 73.58 74.10 79.71 76.35 76.09
Gemini-2.5-Flash 72.43 73.34 70.53 78.32 74.92 75.00
GPT4o 77.67 78.23 73.40 80.07 79.73 77.95
VideoLLaMA3-7B 60.43 62.05 62.07 66.97 61.41 62.70
VideoLLaMA3-7B + UVLM 70.52 71.39 70.17 76.41 73.39 73.03
Claude-3.7-Sonnet Claude-3.7-Sonnet 76.82 73.90 83.70 82.91 75.55 78.99
Gemini-2.5-Flash 79.81 76.44 83.39 83.89 78.15 80.74
GPT4o 71.78 68.41 65.94 73.10 70.67 71.92
VideoLLaMA3-7B 58.65 52.93 60.26 63.12 54.06 59.5
VideoLLaMA3-7B + UVLM 69.23 65.31 70.79 74.57 67.93 71.51
Gemini-2.5-Flash Claude-3.7-Sonnet 81.54 80.85 83.65 90.48 79.97 82.20
Gemini-2.5-Flash 83.57 82.19 82.49 91.55 83.29 83.52
GPT4o 76.60 73.00 65.10 80.69 73.47 73.85
VideoLLaMA3-7B 46.93 44.21 50.98 51.84 46.90 49.59
VideoLLaMA3-7B + UVLM 60.84 56.54 60.36 71.42 60.77 65.57
We report overall performance on these benchmark in Table 4. Qualitative comparisons are shown in Figure 5. Specifically, UVLM significantly enhances the model’s capabilities in spatial reasoning and temporal understanding, enabling more precise and fine-grained comprehension across diverse video scenarios. After fine-tuning with our dataset, the model demonstrates stronger generalization in recognizing object orientation, interpreting target actions, and performing inference. For example, in the first case shown in Figure  5, the question ’What unusual act does the man perform in the video?’ requires accurate judgment of the object’s spatial orientation. The baseline model incorrectly responds with ’He places the broom horizontally on the wall’, failing to capture the actual positioning of the object. In contrast, the model fine-tuned on our dataset correctly identifies the broom as being placed vertically and answers ’He places the broom vertically on the wall’, demonstrating a more acute perception of spatial structure.
Our findings suggest a simple yet effective strategy for building universally capable VLMs: supplement standard web video with challenge domains exhibiting extreme physics and optics. Underwater footage acts as an “immunity booster,” teaching the model to cope with severe domain shifts and improving generalization more efficiently than scaling homogeneous data. Future work will explore other out-of-distribution sources such as infrared, aerial drone, and egocentric streams.
Appendix G F. Evaluation with Different Judgment Models
We conduct a comparative analysis of evaluation using three foundational judges: GPT-4o, Claude-3.7-Sonnet, and Gemini-2.5-Flash. By comparing the resulting scores, we seek to verify that UVLM-induced improvements are robust to evaluator choice and to identify any rating bias or self-preference each judge may exhibit. These findings will validate the generality of the fine-tuning effect and guide future work on selecting, weighting, or calibrating evaluation protocols, ensuring fairer and more credible conclusions.
As shown in Table 5, integrating UVLM produces clear performance gains across all six subjective evaluation dimensions under every judge, with each judge exhibiting its own preference pattern.
The smallest improvement appears under GPT-4o, where VideoLLaMA3-7B’s Overall score rises from 62.70 to 73.03, a gain of +10.33. Claude delivers the most balanced assessments, assigning Overall scores that lie between those of the other two judges. Gemini reports the largest boost: the Overall score climbs from 49.59 to 65.57, an increase of +15.98. In every case, the introduction of UVLM raises the Overall score by at least ten points, demonstrating robust and consistent benefits across diverse judges and model settings.
Notably, Gemini tends to award higher ratings when it evaluates its own generations: most strikingly, 82.19 for Visual Perception Accuracy, 91.55 for Environmental Description Accuracy, 83.29 for Species Behavior Matching, and a lofty Overall score of 83.52. This phenomenon suggests a degree of self-preference bias in Gemini, particularly regarding recognition of visual detail and ecological perception ability.
Appendix H G. Evaluating How Underwater Image Quality Influences Model Accuracy
Figure 6: Image quality influence.
To explore the effects of underwater image quality degradation on video perception, Figure 6 presents the image quality degradation curve along with several underwater video perception metrics. Here, we employ NIQE to assess image quality degradation, where a higher NIQE value indicates more severe degradation. Specifically, we randomly selected nine samples. The NIQE score was computed on ten randomly selected frames, and the average of the ten frames was taken as the final NIQE value. For easy comparison, we ranked the videos in descending order of NIQE.
Experimental results demonstrate a clear correlation between improved image quality and enhanced perception performance. As the NIQE scores decreased from approximately 16 (indicating poorer quality) to 12 (better quality), we observed progressive improvements across all evaluation metrics. The overall accuracy increased by approximately 20%, while semantic accuracy showed particularly significant gains, rising from 70 to 90. Visual perception accuracy similarly exhibited an upward trend with quality improvement.
These findings substantiate the necessity of considering underwater-specific conditions, highlighting the importance of establishing a specialized benchmark for underwater video perception tasks